# When Do Neural Networks Learn World Models?

### v4

@@ E.1 Technical Lemmas
Proof of (i). If $U$ is compatible, then by Lemma E.14 we have $\mathrm{deg}_{U}(f)=\mathrm{deg}(f)$ for any Boolean function $f$. This immediately gives the equivalence between (71) and (56) and hence their minimizers. We thus have that $\Phi^{*}$ learns the world model up to negations and permutations as in Theorem 4.8.
Proof of (ii). If $U$ is not compatible, then due to the invertibility of $U$, there exists at least one $h^{\prime\prime}\in{\mathcal{F}}^{d}_{k}$ such that $\mathrm{deg}_{U}(h^{\prime\prime}\circ T^{-1})>\mathrm{deg}(h^{\prime\prime})$ for some $k<d$. Since composing Boolean functions with degree-$1$ transforms does not change the degree of functions, we have $\mathrm{deg}_{U}(h^{\prime\prime}\circ T^{-1})>\mathrm{deg}(h^{\prime\prime})$ and hence $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})>\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime})$.
@@
Recall that $\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})=\mathrm{deg}(U^{-1}(h^{\prime}\circ T^{-1}))$. Therefore, to ensure that $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})=\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime})$ for $k=1$, we thus must have $T^{-1}_{j}=\chi_{S}$ for some $S\subseteq[d]$ and $j\in[d]$, with $|S|=k$. This means
@@
which gives the desired result. ∎
@@ F.1 Polynomial Extrapolation
We consider fitting and extrapolating degree-$n$ polynomials with the form $P_{n}(x)=\sum_{i=0}^{n}a_{i}x^{n}$. Given an input $x\in{\mathbb{R}}$, the label is given by $y=P_{n}(x)$. In our experiments, we consider three families of polynomials with degree $1$, $2$, and $3$. In each family, every coefficient $a_{i},i\in\{0,1,\ldots,n\}$ is uniformly sampled from $[0,1)$. In our experiments, we sample $50$ polynomials in each family for the violin plots. Other data parameters are as follows:
  • Training, validation, and test data are uniformly sampled from $[-1,1)$, $[-1,1)$, and $[-2,2)$, respectively.
@@ F.2 Learning Physical Laws
  • Initial learning rate is randomly sampled from $[1e-6,1e-3]$. We use a cosine learning rate scheduler.
  • Weight decay is set to $41e-$ 4.
@@ G.1 Polynomial Extrapolation
In the main text, we report extrapolation results on degree-$2$ and degree-$3$ polynomials in Figure 2(b); for completeness, here we also report extrapolation results on degree-$1$ polynomials, i.e., linear functions. As shown by (Xu et al., 2021), ReLU MLPs can also extrapolate well in this setting. The violin plots of the test MSE of both the ReLU MLP and our model are in Figure 8. We can see that while both models achieve a much smaller extrapolation error compared to those in extrapolating higher degree polynomials, our model still outperforms the ReLU MLP. We also provide more examples for extrapolating degree-$1$, degree-$2$, and degree-$3$ polynomials in Figure 9, Figure 10, and Figure 11, respectively.
@@ G.2 Learning Physical Laws
Figure 9: Selected examples for degree-$1$ polynomial extrapolation. Shaded regions indicate training regions.
Figure 10: Selected examples for degree-$2$ polynomial extrapolation. Shaded regions indicate training regions.
Figure 11: Selected examples for degree-$3$ polynomial extrapolation. Shaded regions indicate training regions.

### v5

@@ E.1 Technical Lemmas
Proof of (i). If $U$ is compatible, then by Lemma E.14 we have $\mathrm{deg}_{U}(f)=\mathrm{deg}(f)$ for every Boolean function $f$. This immediately gives the equivalence between (71) and (56) and hence their minimizers. We thus conclude that $\Phi^{*}$ learns the world model up to negations and permutations (i.e., degree-$1$ transforms) as in Theorem 4.8.
Proof of (ii). By the proof of Theorem 4.8 (Section E.6), $\Phi^{*}$ learns the world model up to degree-$1$ transforms if and only if $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})=\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime})$ holds for every $k\in[d]$, where $T=\Phi^{*}\circ\psi$ with $\mathrm{deg}(T_{j})=1,\,\forall j\in[n]$. If $U$ is not compatible, then there must exist $h^{\prime\prime}\in{\mathcal{F}}^{d}_{k}$ such that $\mathrm{deg}_{U}(h^{\prime\prime})>\mathrm{deg}(h^{\prime\prime})$ for some $k<d$. Since composing Boolean functions with degree-$1$ transforms does not change the degree of functions, we have $\mathrm{deg}_{U}(h^{\prime\prime}\circ T^{-1})>\mathrm{deg}(h^{\prime\prime})$ and hence $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})>\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime})$ for every degree-$1$ transform $T$.
@@
Recall that $\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})=\mathrm{deg}(U^{-1}(h^{\prime}\circ T^{-1}))$. To ensure that $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}_{U}(h^{\prime}\circ T^{-1})=\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime})$ for $k=1$, we must have $T^{-1}_{j}=\chi_{S}$ for some $S\subseteq[d]$ and $j\in[d]$ with $|S|=k$. This gives
@@
which completes the proof. ∎
@@ F.1 Polynomial Extrapolation
We consider fitting and extrapolating degree-$n$ polynomials with the form $P_{n}(x)=\sum_{i=0}^{n}a_{i}x^{n}$. Given an input $x\in{\mathbb{R}}$, the label is given by $y=P_{n}(x)$. In our experiments, we consider three families of polynomials with degree $1$, 22, and $3$. In each family, every coefficient $a_{i},i\in\{0,1,\ldots,n\}$ is uniformly sampled from $[0,1)$. In our experiments, we sample $50$ polynomials in each family for the violin plots. Other data parameters are as follows:
  • Training, validation, and test data are uniformly sampled from [1,1)[-1,1), [1,1)[-1,1), and [2,2)[-2,2), respectively.
@@ F.2 Learning Physical Laws
  • Initial learning rate is randomly sampled from [1e6,1e3][1e-6,1e-3]. We use a cosine learning rate scheduler.
  • Weight decay is set to 1e41e-4.
@@ G.1 Polynomial Extrapolation
In the main text, we report extrapolation results on degree-$2$ and degree-$3$ polynomials in Figure 2(b); for completeness, here we also report extrapolation results on degree-$1$ polynomials, i.e., linear functions. As shown by (Xu et al., 2021), ReLU MLPs can also extrapolate well in this setting. The violin plots of the test MSE of both the ReLU MLP and our model are in Figure 8. We can see that while both models achieve a much smaller extrapolation error compared to those in extrapolating higher degree polynomials, our model still outperforms the ReLU MLP. We also provide more examples for extrapolating degree-$1$, degree-22, and degree-$3$ polynomials in Figure 9, Figure 10, and Figure 11, respectively.
@@ G.2 Learning Physical Laws
Figure 9: Selected examples for degree-11 polynomial extrapolation. Shaded regions indicate training regions.
Figure 10: Selected examples for degree-22 polynomial extrapolation. Shaded regions indicate training regions.
Figure 11: Selected examples for degree-33 polynomial extrapolation. Shaded regions indicate training regions.